# On the non-archimedean Monge–Ampère equation in mixed characteristic

### v2

@@ 1. Introduction
Yuan and Zhang proved uniqueness up to scaling for solutions of the non-archimedean Monge–Ampère equation [YZ17]. In a groundbreaking work, Boucksom, Favre and Jonsson solved the non-archimedean Monge–Ampère equation over a complete discretely valued field $K$ of residue characteristic zero assuming that $X$ is smooth and that the support of the Radon measure $\mu$ is contained in the skeleton of an SNC-model (a projective regular model with special fiber having simple normal crossing support) [BFJ15]. They assumed also that $K$ is a completion of the function field of a curve at a closed point. This geometric condition was later removed by Burgos, Jell, Martin and the last two authors of this paper [${}^{+}$20]. Boucksom, Favre, and Jonsson used a variational approach to solve (1.1) which relies crucially on the continuity of the semipositive envelope of a continuous metric of $L$. They show continuity of this envelope by using multiplier ideals on SNC-models [BFJ16]. In their arguments, Hironaka’s resolution of singularities plays an important role in order to have sufficiently many SNC-models at hand. Apart from the multiplier ideals, it is precisely here where residue characteristic zero is used again.
@@
In this paper, we deal with the Monge–Ampère problem for arbitrary smooth projective varieties over a complete discretely valued field $K$ of mixed characteristic $(0,p)$. For a complete local noetherian domain $R$ of mixed characteristic, a theory of test ideals was introduced by Ma and Schwede based on perfectoid ideas [MS18, MS21]. Using perfectoid methods, prismatic techniques, and a $p$-adic Riemann-Hilbert correspondence, Bhatt was able in 2020 to show a variant of Kodaira vanishing ‘up to finite covers’ in mixed characteristic [Bha20]. Applications to the minimal model program were given by Bhatt, Ma, Patakfalvi, Schwede, Tucker, Waldron, and Witaszek [${}^{+}$21] and independently by Takamatsu and Yoshikawa [TY21]. For projective normal schemes over $R$, these ideas were extended by Hacon, Lamarche and Schwede. They introduced $+$-test ideals and showed global generation results [HLS21]. Their construction, which replaces spaces of global sections by so-called spaces of $+$-stable sections ${}^{+}$21, TY21], is recalled in §3. Hacon, Lamarche and Schwede conjecture that subadditivity holds for their $+$-test ideals [HLS21, Conjecture 8.3]. Observe that subaddivity is well known for multiplier resp. test ideals in the equicharacteristic case. While this conjecture remains open, a modified version of $+$-test ideals, which we call perturbation friendly global test ideals in this article, was introduced by Bhatt, Ma, Patakfalvi, Schwede, Tucker, Waldron, and Witaszek [${}^{+}$24], again benefiting from a $p$-adic Riemann-Hilbert correspondence. These perturbation friendly test ideals enjoy properties as nice as the $+$-test ideals of Hacon, Lamarche and Schwede, and in addition satisfy the subadditivity property. We refer to §4 for details.
The contribution of this paper is to show that global $+$-test ideals and perturbation friendly global test ideals allow applications to the non-archimedean Monge–Ampère problem in mixed characteristic. More precisely, we prove the following results.
@@ 2. Model metrics, curvature forms and psh envelopes
Let $X$ be a proper variety over a non-archimedean field $K$. A model of $X$ is a proper flat scheme ${\mathscr{X}}$ over $S\coloneqq\mathrm{Spec}~{}K^{\circ}$ together with an isomorphism $h$ from $X$ to the generic fiber ${\mathscr{X}}_{\eta}$ of the $S$-scheme ${\mathscr{X}}$. The special fiber of ${\mathscr{X}}$ is denoted by ${\mathscr{X}}_{s}$. Let $L$ be a line bundle over $X$. A model of $(X,L)$ is given by a model ${\mathscr{X}}$ of $X$ and a line bundle ${\mathscr{L}}$ over ${\mathscr{X}}$ with an isomorphism of $L$ to $h^{*}({\mathscr{L}}|_{{\mathscr{X}}_{\eta}})$.
[…]
@@
Given a model ${\mathscr{X}}$ of $X$ one defines $N^{1}({\mathscr{X}}/S)_{{\mathbb{Q}}}$ as the quotient of $\mathrm{Pic}({\mathscr{X}})_{{\mathbb{Q}}}$ by the subspace generated by line bundles whose restriction to every closed curve $C$ in the special fiber ${\mathscr{X}}_{s}$ has degree zero. Call $\alpha\in N^{1}({\mathscr{X}}/S)\coloneqq N^{1}({\mathscr{X}}/S)_{\mathbb{Q}}\otimes_{{\mathbb{Q}}}{\mathbb{R}}$ nef if $\alpha\cdot C\geq 0$ for all such curves.
@@
where the direct limit is taken over all isomorphism classes of models of $X$. For the model metric induced by a model ${\mathscr{L}}$ of $L^{\otimes m}$, its curvature form $c_{1}(L,\|\ \|_{{\mathscr{L}}})$ is the image of ${\mathscr{L}}^{\otimes\frac{1}{m}}\in N^{1}({\mathscr{X}}/S)_{{\mathbb{Q}}}$ in ${\mathcal{Z}}^{1,1}(X)$. By construction we have a map $dd^{c}\colon{\mathscr{D}}(X)\to{\mathcal{Z}}^{1,1}(X)$ given by $g\mapsto c_{1}({\mathcal{O}}_{X},\|\ \|\cdot\mathrm{e}^{-g})$. Call a closed $(1,1)$-form $\theta\in{\mathcal{Z}}^{1,1}(X)$ semipositive if it can be represented by a nef element in $N^{1}({\mathscr{X}}/S)$ for some model ${\mathscr{X}}$. Denote by $N^{1}(X)$ the quotient of $\mathrm{Pic}(X)_{\mathbb{R}}$ by numerical equivalence. The map {}$\{\phantom{a}\}\colon{\mathcal{Z}}^{1,1}(X)\to N^{1}(X)$ induced by the restriction maps $N^{1}({\mathscr{X}}/S)\to N^{1}(X)$ sends a closed $(1,1)$-form $\theta$ to its de Rham class $\{\theta\}$. A class in $N^{1}(X)$ is called ample if it is an ${\mathbb{R}}_{>0}$-linear combination of classes induced by ample line bundles on $X$.
@@
on $X^{\rm an}$ where $V$ runs through the irreducible components of ${\mathscr{X}}_{s}$, $x_{V}\in X^{{\rm an}}$ is the unique preimage of the generic point $\xi_{V}$ of $V$ under $\mathrm{red}_{\mathscr{X}}$, and $\delta_{x_{V}}$ denotes the Dirac probability measure supported in $x_{V}$. For generalization of the Monge–Ampère measure to all $\theta\in{\mathcal{Z}}^{1,1}(X)$ with ample de Rham class $\{\theta\}$ and to more general classes of $\theta$-psh functions, we refer to [BFJ15], [GJKM19] and Section 8.
@@ 3. Global test ideals in mixed characteristic
In this section, we introduce the global $+$-test ideals defined and studied by Hacon, Lamarche and Schwede in [HLS21]. The $+$-test ideals form a mixed characteristic analogue of the theories of multiplier ideals in characteristic zero and test ideals in positive characteristic. The theory is based on the notion of $+$-stable sections introduced in [${}^{+}$21, TY21].
@@ 3.1. The test ideals of Hacon–Lamarche–Schwede
For a reflexive sheaf ${\mathscr{M}}={\mathcal{O}}_{\mathscr{X}}(M)$ associated with a divisor $M$ and an effective ${\mathbb{Q}}$-divisor $B$ on ${\mathscr{X}}$, the subspace of $+$-stable sections of the adjoint line bundle $\omega_{\mathscr{X}}\otimes{\mathscr{M}}$ relative to $B$
@@
If $M-B$ is ${\mathbb{Q}}$-Cartier, then Definition 3.1 holds as well if $f$ runs through the alterations $f\colon{\mathscr{Y}}\to{\mathscr{X}}$ from a normal integral scheme ${\mathscr{Y}}$ together with an embedding $\kappa({\mathscr{Y}})\hookrightarrow\overline{\kappa({\mathscr{X}})}$ [HLS21, Lemma 3.8].
@@
The sequence $({\mathscr{J}}_{i})_{i\in{\mathbb{N}}}$ is increasing and becomes stationary. Define the $+$-test submodule $\tau_{+}(\omega_{\mathscr{X}},B)$ to be ${\mathscr{J}}_{i}$ for $i\gg 0$ [HLS21, Definition 4.3].
@@ 4. Perturbation friendly global test ideals
In this section, we present the perturbation friendly global test ideals after Bhatt, Ma, Patakfalvi, Schwede, Tucker, Waldron, and Witaszek [${}^{+}$24]. They are suitable for our purpose as subadditivity is known for them in contrast to the test ideals considered in the previous section.
@@
Let $B$ be a ${\mathbb{Q}}$-divisor on ${\mathscr{X}}$. By [${}^{+}$24, Proposition 7.14(f)], there is a unique perturbation friendly test module $\tau(\omega_{\mathscr{X}},B)$ such that there exists an effective Cartier divisor $G^{0}$ with the property that
@@
Both the test module $\tau(\omega_{\mathscr{X}},B)$ and the test ideal $\tau({\mathcal{O}}_{\mathscr{X}},B)$ are coherent [${}^{+}$24, Definition 7.12, Theorem 7.13]. If $B$ is effective, then the fractional ideal sheaf $\tau({\mathcal{O}}_{\mathscr{X}},B)$ is indeed an ideal sheaf in ${\mathcal{O}}_{{\mathscr{X}}}$ [${}^{+}$24, Definition 7.18].
@@
Remark 4.2.
@@
(i) By [${}^{+}$24, Corollary 7.19(b)], there exists a Cartier divisor $G^{0}$ with the property that $\tau({\mathcal{O}}_{\mathscr{X}},B)=\tau_{+}({\mathcal{O}}_{\mathscr{X}},B+\varepsilon G)$ for all divisors $G\geq G^{0}$ on ${\mathscr{X}}$ and all $0<\varepsilon\ll 1$ (depending on $G$).
(ii) If $B$ is a ${\mathbb{Q}}$-divisor and $F$ is a divisor on ${\mathscr{X}}$, then (3.3), (4.1), and (4.2) yield
@@
(4.4) $\tau({\mathcal{O}}_{\mathscr{X}},B+F)=\tau({\mathcal{O}}_{\mathscr{X}},B)\otimes{\mathcal{O}}_{\mathscr{X}}(-F)$
(iii) (Subadditivity) The perturbation friendly test ideals satisfy the subadditivity property [${}^{+}$24, Theorem 7.20(e)], i.e. we have
@@
(iv) (Effective global generation) Let $B$ be effective and let $D$ be a divisor such that the ${\mathbb{Q}}$-divisor $D-K_{\mathscr{X}}-B$ is big and nef, and let $H$ be a globally generated ample divisor. Then $\tau({\mathcal{O}}_{\mathscr{X}},B)\otimes{\mathcal{O}}_{\mathscr{X}}(nH+D)$ is globally generated by $\mathbf{B}^{0}({\mathscr{X}},B,{\mathcal{O}}_{\mathscr{X}}(nH+D))$ for all $n\geq\mathrm{dim}({\mathscr{X}}\otimes_{R}k)$ [${}^{+}$24, Corollary 7.22].
(v) ($\tau({\mathcal{O}}_{\mathscr{X}},0)$ is vertical) Recall that a coherent ideal of ${\mathscr{X}}$ is called vertical if its support is contained in the special fiber. Let $X\coloneqq{\mathscr{X}}\otimes_{R}K$ denote the generic fibre of ${\mathscr{X}}$, put $K_{X}\coloneqq K_{\mathscr{X}}|_{X}$, and recall that the Grauert–Riemenschneider sheaf $\mathscr{J}(X,\omega_{X})$ on the regular scheme $X$ equals $\omega_{X}$ [${}^{+}$24, Definition A.1]. We conclude from [${}^{+}$24, Proposition 7.14(c)], and formula (4.3) above that
@@
It follows that the ideal sheaf $\tau({\mathcal{O}}_{\mathscr{X}},0)$ is vertical.
@@
Let $D$ be a $\mathbb{Q}$-divisor with Itaka dimension $\kappa(D)\geq 0$, so $mD$ is a divisor such that $|mD|\neq\emptyset$ for all large and sufficiently divisible $m\in\mathbb{N}$, and $\lambda\in{\mathbb{Q}}_{>0}$. Define the asymptotic perturbation friendly test ideal of $|D|$ by
(4.7) $\tau({\mathcal{O}}_{\mathscr{X}},\lambda\cdot||D||)\coloneqq\bigcup_{m>0}\tau\Bigl{(}{\mathcal{O}}_{\mathscr{X}},\frac{\lambda}{m}\cdot|mD|\Bigr{)}.$
@@
    (4.8) $\tau({\mathcal{O}}_{\mathscr{X}},\lambda\cdot||D||)=\tau\Bigl{(}{\mathcal{O}}_{\mathscr{X}},\frac{\lambda}{m}\cdot|mD|\Bigr{)}.$
@@
    If $|D|\neq\emptyset$, then $\tau({\mathcal{O}}_{\mathscr{X}},\lambda\cdot|D|)\subset\tau({\mathcal{O}}_{\mathscr{X}},\lambda\cdot||D||)$.
@@
    If $\lambda<\mu$, $(\lambda,\mu)\in{\mathbb{Q}}_{>0}^{2}$, then $\tau({\mathcal{O}}_{\mathscr{X}},\mu\cdot||D||)\subset\tau({\mathcal{O}}_{\mathscr{X}},\lambda\cdot||D||)$.
@@
    If $k\in\mathbb{N}_{>0}$, then $\tau({\mathcal{O}}_{\mathscr{X}},\lambda\cdot||D||)=\tau({\mathcal{O}}_{\mathscr{X}},\frac{\lambda}{k}\cdot||kD||)$.
@@
Let $D$ be a $\mathbb{Q}$-divisor on ${\mathscr{X}}$ with Itaka dimension $\kappa(D)\geq 0$. (i) Let ${\mathscr{H}}={\mathcal{O}}_{\mathscr{X}}(H)$ be a globally generated ample line bundle, $E$ be a divisor, and $\lambda\in{\mathbb{Q}}_{>0}$. Let $n=\dim{\mathscr{X}}\otimes_{R}k$. If $E-K_{\mathscr{X}}-\lambda D$ is big and nef, then
(4.9) $\tau({\mathcal{O}}_{\mathscr{X}},\lambda\cdot||D||)\otimes{\mathcal{O}}_{\mathscr{X}}(nH+E)$
@@
(i) follows from [${}^{+}$24, Remark 8.36]. For the convenience of the reader we give a proof following [HLS21, Lemma 7.5 (e)]. For sufficiently divisible $m\in{\mathbb{N}}$ we conclude from (4.8) that
$\tau({\mathcal{O}}_{\mathscr{X}},\lambda\cdot||D||)\otimes{\mathcal{O}}_{\mathscr{X}}(nH+E)$ $=\tau\Bigl{(}{\mathscr{O}}_{\mathscr{X}},\frac{\lambda}{m}\cdot|mD|\Bigr{)}\otimes{\mathcal{O}}_{\mathscr{X}}(nH+E)$
@@
$\tau\bigl{(}{\mathscr{O}}_{\mathscr{X}},q\cdot||D||\bigr{)}=\tau\Bigl{(}{\mathscr{O}}_{\mathscr{X}},\frac{q}{m}\cdot|mD|\Bigr{)}.$
@@
$\tau\bigl{(}{\mathscr{O}}_{\mathscr{X}},q\cdot||D||\bigr{)}\subset\tau\Bigl{(}{\mathscr{O}}_{\mathscr{X}},\frac{1}{m}\cdot|mD|\Bigr{)}^{q}\subset\tau\bigl{(}{\mathcal{O}}_{\mathscr{X}},\|D\|\bigr{)}^{q}$

### v3

@@ 1. Introduction
Yuan and Zhang proved uniqueness up to scaling for solutions of the non-archimedean Monge–Ampère equation [YZ17]. In a groundbreaking work, Boucksom, Favre and Jonsson solved the non-Archimedean Monge–Ampère equation over a complete discretely valued field $K$ of residue characteristic zero assuming that $X$ is smooth and that the support of the Radon measure $\mu$ is contained in the skeleton of an SNC-model (a projective regular model with special fiber having simple normal crossing support) [BFJ15]. They assumed also that $K$ is a completion of the function field of a curve at a closed point. This geometric condition was later removed by Burgos, Jell, Martin and the last two authors of this paper [BGJ+20]. Boucksom, Favre, and Jonsson used a variational approach to solve (1.1) which relies crucially on the continuity of the semipositive envelope of a continuous metric of $L$. They show continuity of this envelope by using multiplier ideals on SNC-models [BFJ16]. In their arguments, Hironaka’s resolution of singularities plays an important role in order to have sufficiently many SNC-models at hand. Apart from the multiplier ideals, it is precisely here where residue characteristic zero is used again.
@@
In this paper, we deal with the Monge–Ampère problem for arbitrary smooth projective varieties over a complete discretely valued field $K$ of mixed characteristic $(0,p)$. For a complete local noetherian domain $R$ of mixed characteristic, a theory of test ideals was introduced by Ma and Schwede based on perfectoid ideas [MS18, MS21]. Using perfectoid methods, prismatic techniques, and a $p$-adic Riemann-Hilbert correspondence, Bhatt was able in 2020 to show a variant of Kodaira vanishing ‘up to finite covers’ in mixed characteristic [Bha20]. Applications to the minimal model program were given by Bhatt, Ma, Patakfalvi, Schwede, Tucker, Waldron, and Witaszek [BMP+21] and independently by Takamatsu and Yoshikawa [TY21]. For projective normal schemes over $R$, these ideas were extended by Hacon, Lamarche and Schwede. They introduced ++-test ideals and showed global generation results [HLS21]. Their construction, which replaces spaces of global sections by so-called spaces of ++-stable sections from[BMP+21, TY21], is recalled in §3. Hacon, Lamarche and Schwede conjecture that subadditivity holds for their ++-test ideals [HLS21, Conjecture 8.3]. Observe that subaddivity is well known for multiplier resp. test ideals in the equicharacteristic case. While this conjecture remains open, a modified version of ++-test ideals, which we call perturbation friendly global test ideals in this article, was introduced by Bhatt, Ma, Patakfalvi, Schwede, Tucker, Waldron, and Witaszek [BMP+24], again benefiting from a $p$-adic Riemann-Hilbert correspondence. These perturbation friendly test ideals enjoy properties as nice as the ++-test ideals of Hacon, Lamarche and Schwede, and in addition satisfy the subadditivity property. We refer to §4 for details.
The contribution of this paper is to show that global ++-test ideals and perturbation friendly global test ideals allow applications to the non-Archimedean Monge–Ampère problem in mixed characteristic. More precisely, we prove the following results.
@@ 2. Model metrics, curvature forms and psh envelopes
Let $X$ be a proper variety over a non-Archimedean field $K$. Recall that the topological space underlying the Berkovich analytification $X^{\mathrm{an}}$ of $X$ consists of pairs $(p,|\phantom{a}|_{p})$ with $p\in X$ and $|\phantom{a}|_{p}$ an absolute value on $\kappa(p)=\mathcal{O}_{X,p}/\mathfrak{m}_{X,p}$ that extends $|\phantom{a}|_{K}$ and is equipped with coarsest topology such that the map $\pi\colon X^{\rm an}\rightarrow X$, $(p,|\phantom{a}|_{p})\mapsto p$ is continuous and for all Zariski open subsets $U$ of $X$ and all regular functions $f\in\mathcal{O}_{X}(U)$ the map $|f|\colon U^{\mathrm{an}}=\pi^{-1}(U)\rightarrow{\mathbb{R}},\,(p,|\phantom{a}|_{p})\mapsto|f(p)|:=|f+\mathfrak{m}_{X,p}|_{p}$ is continuous as well.
A model of $X$ is a proper flat scheme $\mathscr{X}$} over $S\coloneqq\mathrm{Spec}~K^{\circ}$ together with an isomorphism $h$ from $X$ to the generic fiber ${\mathscr{X}}_{\eta}$ of the $S$-scheme $\mathscr{X}$}. The special fiber of $\mathscr{X}$} is denoted by ${\mathscr{X}}_{s}$. Given a model $\mathscr{X}$}, the valuative criterion for properness yields a natural reduction map $\mathrm{red}_{\mathscr{X}}\colon X^{\rm an}\to{\mathscr{X}}_{s}$. Let $L$ be a line bundle over $X$. A model of $(X,L)$ is given by a model $\mathscr{X}$} of $X$ and a line bundle $\mathscr{L}$} over $\mathscr{X}$} with an isomorphism of $L$ to $h^{*}({\mathscr{L}}|_{{\mathscr{X}}_{\eta}})$. A continuous metric $\|\ \|$ of $L$ associates with each section $s\in\Gamma(U,L)$ on some Zariski open subset $U$ of the variety $X$ a continuous function $\|s\|\colon U^{\mathrm{an}}\to[0,\infty)$ such that one has $\|f\cdot s\|=|f|\cdot\|s\|$ for each regular function $f$ in ${\mathcal{O}}_{X}(U)$. It is furthermore required that $\|s\|>0$, if $s$ is a nowhere vanishing section of $L$.
Let $({\mathscr{X}},{\mathscr{L}})$ be a model of $(X,L^{\otimes m})$ for some $m\in{\mathbb{N}}_{>0}$. The model metric $\|\ \|_{{\mathscr{L}}}$ of $L$ over $X^{{\rm an}}$ is determined by $\|s\|_{{\mathscr{L}}}:=\sqrt[m]{|g|}$ on $U^{{\rm an}}\cap\mathrm{red}^{-1}({\mathscr{U}}_{s})$, where $\mathscr{U}$} is open in $\mathscr{X}$}, $s$ is a section of $L$ over $U\coloneqq X\cap{\mathscr{U}}$, $b$ is a nowhere vanishing section of $\mathscr{L}$} over $\mathscr{U}$}, $g\in{\mathcal{O}}_{X}(U)$ is a regular function such that $s^{\otimes m}=gb$ over $U$, and $\mathrm{red}_{\mathscr{X}}\colon X^{\rm an}\to{\mathscr{X}}_{s}$ is the reduction map. A model metric $\|\ \|$ on ${\mathcal{O}}_{X}^{{\rm an}}$ induces a so-called model function $\log\|1\|\colon X^{{\rm an}}\to{\mathbb{R}}$. The $\mathbb{Q}$}-vector space of all model functions on $X$ is denoted by ${\mathscr{D}}(X)$.
@@
Given a model $\mathscr{X}$} of $X$ one defines $N^{1}({\mathscr{X}}/S)_{{\mathbb{Q}}}$ as the quotient of the $\mathbb{Q}$}-vector space $\mathrm{Pic}({\mathscr{X}})_{{\mathbb{Q}}}:=\mathrm{Pic}({\mathscr{X}})\otimes_{\mathbb{Z}}{\mathbb{Q}}$ by the subspace generated by line bundles whose restriction to every closed curve $C$ in the special fiber ${\mathscr{X}}_{s}$ has degree zero. Call $\alpha\in N^{1}({\mathscr{X}}/S)\coloneqq N^{1}({\mathscr{X}}/S)_{\mathbb{Q}}\otimes_{{\mathbb{Q}}}{\mathbb{R}}$ nef if $\alpha\cdot C\geq 0$ for all such curves.
@@
where the direct limit is taken over all isomorphism classes of models of $X$ and the transition maps are induced by pullback between dominating models. For the model metric induced by a model $\mathscr{L}$} of $L^{\otimes m}$, its curvature form $c_{1}(L,\|\ \|_{{\mathscr{L}}})$ is the image of ${\mathscr{L}}^{\otimes\frac{1}{m}}\in N^{1}({\mathscr{X}}/S)_{{\mathbb{Q}}}$ in ${\mathcal{Z}}^{1,1}(X)$. By construction we have a map $dd^{c}\colon{\mathscr{D}}(X)\to{\mathcal{Z}}^{1,1}(X)$ given by $g\mapsto c_{1}({\mathcal{O}}_{X},\|\ \|\cdot\mathrm{e}^{-g})$. Call a closed $(1,1)$-form $\theta\in{\mathcal{Z}}^{1,1}(X)$ semipositive if it can be represented by a nef element in $N^{1}({\mathscr{X}}/S)$ for some model $\mathscr{X}$}. Denote by $N^{1}(X)$ the quotient of $\mathrm{Pic}(X)_{\mathbb{R}}:=\mathrm{Pic}(X)\otimes_{\mathbb{Z}}{\mathbb{R}}$ by numerical equivalence. The map $\{\phantom{a}\}\colon{\mathcal{Z}}^{1,1}(X)\to N^{1}(X)$ induced by the restriction maps $N^{1}({\mathscr{X}}/S)\to N^{1}(X)$ sends a closed $(1,1)$-form $\theta$ to its de Rham class $\{\theta\}$. A class in $N^{1}(X)$ is called ample if it is an ${\mathbb{R}}_{>0}$-linear combination of classes induced by ample line bundles on $X$.
@@
on $X^{\rm an}$ where $V$ runs through the irreducible components of ${\mathscr{X}}_{s}$, $x_{V}\in X^{{\rm an}}$ is the unique preimage of the generic point $\xi_{V}$ of $V$ under the reduction map $\mathrm{red}_{\mathscr{X}}\colon X^{\rm an}\to{\mathscr{X}}_{s}$, and $\delta_{x_{V}}$ denotes the Dirac probability measure supported in $x_{V}$. For generalization of the Monge–Ampère measure to all $\theta\in{\mathcal{Z}}^{1,1}(X)$ with ample de Rham class $\{\theta\}$ and to more general classes of $\theta$-psh functions, we refer to [BFJ15], [GJKM19] and Section 8.
@@ 3. Global test ideals in mixed characteristic
In this section, we introduce the global ++-test ideals defined and studied by Hacon, Lamarche and Schwede in [HLS21]. The ++-test ideals form a mixed characteristic analogue of the theories of multiplier ideals in characteristic zero and test ideals in positive characteristic. The theory is based on the notion of ++-stable sections introduced in [BMP+21, TY21].
@@
For a reflexive sheaf ${\mathscr{M}}={\mathcal{O}}_{\mathscr{X}}(M)$ associated with a divisor $M$ and an effective $\mathbb{Q}$}-divisor $B$ on $\mathscr{X}$}, the subspace of ++-stable sections of the adjoint line bundle $\omega_{\mathscr{X}}\otimes{\mathscr{M}}$ relative to $B$
@@
If MBM-B is $\mathbb{Q}$}-Cartier, then Definition 3.1 holds as well if $f$ runs through the alterations $f\colon{\mathscr{Y}}\to{\mathscr{X}}$ from a normal integral scheme $\mathscr{Y}$} together with an embedding $\kappa({\mathscr{Y}})\hookrightarrow\overline{\kappa({\mathscr{X}})}$ [HLS21, Lemma 3.8].
@@
The sequence $({\mathscr{J}}_{i})_{i\in{\mathbb{N}}}$ is increasing and becomes stationary. Define the ++-test submodule $\tau_{+}(\omega_{\mathscr{X}},B)$ to be ${\mathscr{J}}_{i}$ for $i\gg 0$ [HLS21, Definition 4.3].
@@ 4. Perturbation friendly global test ideals
In this section, we present the perturbation friendly global test ideals after Bhatt, Ma, Patakfalvi, Schwede, Tucker, Waldron, and Witaszek [BMP+24]. They are suitable for our purpose as subadditivity is known for them in contrast to the test ideals considered in the previous section.
@@
Let $B$ be a $\mathbb{Q}$}-divisor on $\mathscr{X}$}. By [BMP+24, Proposition 7.14(f)], there is a unique perturbation friendly test module $\tau(\omega_{\mathscr{X}},B)$ such that there exists an effective Cartier divisor $G^{0}$ with the property that
@@
Both the test module $\tau(\omega_{\mathscr{X}},B)$ and the test ideal $\tau({\mathcal{O}}_{\mathscr{X}},B)$ are coherent [BMP+24, Definition 7.12, Theorem 7.13]. If $B$ is effective, then the fractional ideal sheaf $\tau({\mathcal{O}}_{\mathscr{X}},B)$ is indeed an ideal sheaf in ${\mathcal{O}}_{{\mathscr{X}}}$ [BMP+24, Definition 7.18].
@@
Theorem 4.2 (Bhatt, Ma, Patakfalvi, Schwede, Tucker, Waldron, Witaszek).
@@
(i) There exists a Cartier divisor $G^{0}$ on $\mathscr{X}$} with the property that
$\tau({\mathcal{O}}_{\mathscr{X}},B)=\tau_{+}({\mathcal{O}}_{\mathscr{X}},B+\varepsilon G)$
(ii) If $B$ is a $\mathbb{Q}$}-divisor and $F$ is a divisor on $\mathscr{X}$}, then
@@
(4.4) $\tau({\mathcal{O}}_{\mathscr{X}},B+F)=\tau({\mathcal{O}}_{\mathscr{X}},B)\otimes{\mathcal{O}}_{\mathscr{X}}(-F).$
(iii) (Subadditivity) The perturbation friendly test ideals satisfy the subadditivity property, i.e. we have
@@
(iv) (Effective global generation) Let $B$ be effective and let $D$ be a divisor such that the $\mathbb{Q}$}-divisor $D-K_{\mathscr{X}}-B$ is big and nef, and let $H$ be a globally generated ample divisor. Then $\tau({\mathcal{O}}_{\mathscr{X}},B)\otimes{\mathcal{O}}_{\mathscr{X}}(nH+D)$ is globally generated by $\mathbf{B}^{0}({\mathscr{X}},B,{\mathcal{O}}_{\mathscr{X}}(nH+D))$ for all $n\geq\mathrm{dim}({\mathscr{X}}\otimes_{R}k)$.
(v) ($\tau({\mathcal{O}}_{\mathscr{X}},0)$ is vertical) The ideal sheaf $\tau({\mathcal{O}}_{\mathscr{X}},0)$ is vertical, i.e. its support is contained in the special fiber.
Proof.
Everything is shown in [BMP+24]. We give precise references. Property (i) holds by [BMP+24, Corollary 7.19(b)] and (ii) follows from (3.3), (4.1), and (4.2). Statement (iii) is the subadditivity property [BMP+24, Theorem 7.20(e)] and (iv) is the effective global generation property [BMP+24, Corollary 7.22]. Recall for (v) that a coherent ideal of $\mathscr{X}$} is called vertical if its support is contained in the special fiber. Let $X\coloneqq{\mathscr{X}}\otimes_{R}K$ denote the generic fibre of $\mathscr{X}$}, put $K_{X}\coloneqq K_{\mathscr{X}}|_{X}$, and recall that the Grauert–Riemenschneider sheaf $\mathscr{J}(X,\omega_{X})$ on the regular scheme $X$ equals $\omega_{X}$ [BMP+24, Definition A.1]. We conclude from [BMP+24, Proposition 7.14(c)], and formula (4.3) above that
@@
It follows that the ideal sheaf $\tau({\mathcal{O}}_{\mathscr{X}},0)$ is vertical. ∎
@@
Let $D$ be a $\mathbb{Q}$-divisor with Iitaka dimension $\kappa(D)\geq 0$, so $mD$ is a divisor such that $|mD|\neq\emptyset$ for all large and sufficiently divisible $m\in\mathbb{N}$, and $\lambda\in{\mathbb{Q}}_{>0}$. Define the asymptotic perturbation friendly test ideal of $|D|$ by
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Let $D$ be a $\mathbb{Q}$-divisor on $\mathscr{X}$} with Iitaka dimension $\kappa(D)\geq 0$.
(i) Let ${\mathscr{H}}={\mathcal{O}}_{\mathscr{X}}(H)$ be a globally generated ample line bundle, $E$ be a divisor, and $\lambda\in{\mathbb{Q}}_{>0}$. Let $n=\dim{\mathscr{X}}\otimes_{R}k$. If $E-K_{\mathscr{X}}-\lambda D$ is big and nef, then
[…]
@@
(i) follows from [BMP+24, Remark 8.36]. For the convenience of the reader we give a proof following [HLS21, Lemma 7.5 (e)]. For sufficiently divisible $m\in{\mathbb{N}}$ we conclude from (4.8) that
[…]
@@
[…]
@@
[…]